OpenStudy (anonymous):

Simplify the expression. 3 + 2 • (8 – 3) A. 16 B. 13 C. 37 D. 25

OpenStudy (anonymous):

D. 25

OpenStudy (anonymous):

how

OpenStudy (anonymous):



OpenStudy (anonymous):

u cute

OpenStudy (anonymous):

3 plus 2 = 5 and 8 minus 3 = 5 and 5x5 = 25

OpenStudy (anonymous):

ok thx

OpenStudy (cwrw238):

pedmas gives order of priority parenthese- exponents-divide-multiply=add-subtract

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

no its 24

OpenStudy (littlebird):

PEMDAS Please excuse my dear aunt Sally Parantheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

PEMDAS Please excuse my dear aunt Sally Parantheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

its 13

OpenStudy (anonymous):

how

OpenStudy (anonymous):

is it 13

OpenStudy (anonymous):

its b

OpenStudy (cwrw238):

correct is 13

OpenStudy (anonymous):

so 13

OpenStudy (anonymous):

yah its 13

OpenStudy (cwrw238):

= 3 + 2 * 5 = 3 + 10 = 13

OpenStudy (anonymous):

not 25

OpenStudy (anonymous):

Are you familiar with PEMDAS? the answer is 13.

OpenStudy (anonymous):

not 25

OpenStudy (anonymous):

@nikkikitty12

OpenStudy (anonymous):